OpenStudy (empty):

I'm trying to find the most general bounds and functions work for this integral:

OpenStudy (empty):

\[\int\limits_a^b \sqrt{ \sum_{n=0}^\infty \left( f(x) \right)^n } dx\]

OpenStudy (empty):

I guess this is kind of silly, f(x) could be anything as long as the value inside the square root is greater than or equal to zero. Haha hmm.

OpenStudy (anonymous):

One thing you can say is that, at least over the interval \([a,b]\), you have to have \(|f(x)|<1\).

OpenStudy (empty):

(changed this function to g(x) to avoid confusion)Really this comes from trying to solve a differential equation of this form: \[y^{(n)} = \tfrac{1}{2} g'(y^{(n-2)})\]

OpenStudy (anonymous):

This looks like something that could be handled or at the very least examined with calculus of variations ... assuming the question can be posed appropriately for it.

OpenStudy (empty):

That's a good idea I don't know how I'd do that, I really am just playing around with this idea in a more general form: \[y'' = \frac{dy'}{dx} = \frac{dy'}{dy}\frac{dy}{dx} = \frac{dy'}{dy}y' \] So if you have something of this form: \[y''=f'(y)\] you can make it separable: \[y' \frac{dy'}{dy} = f'(y)\]

OpenStudy (anonymous):

Provided that \(|f(x)|<1\) for \(a\le x\le b\), \[\int_a^b\sqrt{\sum_{n\ge0}f(x)^n}\,\mathrm{d}x=\int_a^b \sqrt{\frac{1}{1-f(x)}}\,\mathrm{d}x\](along with the usual conditions needed for the integral to exist). Not sure what else can be done with this in this general form.

OpenStudy (empty):

hmmm I wonder if I approximate the sum as for \(f(x)<1\) \[\sum_{n=0}^\infty f(x)^n \approx \int_0^\infty f(x)^ndn = -\ln f(x)\] \[\int_a^b \sqrt{-\ln f(x) } dx\] Basically I'm just trying to find a closed form even if I gotta fudge it now haha.